OpenStudy (anonymous):

Yuri makes his own citrus vinaigrette salad dressing. He likes the dressing best if it contains 30% fruit juice. He has a bottle of dressing mixture that contains 60% juice and 3 cups of dressing that contains 10% juice. How much of the dressing mix with 60% juice should Yuri add to the dressing mix with 10% juice to get the dressing he prefers with 30% juice? Show your work.

pooja195 (pooja195):

60a + 10(1-a) = 1*30 50a = 20 a = 0.4 and 1-a = 0.6. They should be mixed in the ratio 4:6